# PoreTrack3D: A Benchmark for Dynamic 3D Gaussian Splatting in Pore-Scale Facial Trajectory Tracking

### v1

@@ III-F PoreTrack3D Dataset
Based on these calibrated data, we reconstruct a 3D point cloud for the initial frame and generate dense temporal trajectories through our proposed four-stage trajectory reconstruction framework. In total, PoreTrack3D contains over 440,000 trajectories, among which more than 52,000 extend beyond 10 frames. Of these, 68 primary trajectories span the full 150 frames and have been manually reviewed, as illustrated in Fig. 5. These trajectories cover both conventional facial landmarks and pore-scale surface keypoints, enabling quantitative analysis of fine-grained facial deformation. Specifically, among the 68 full-length trajectories, 20 trajectories correspond to facial landmarks and 48 trajectories correspond to pore-scale keypoints. Notably, trajectories longer than 10 frames constitute less than 15% of the total, underscoring the challenge of achieving temporally coherent keypoint tracking under realistic motion and illumination variations.
Fig. 6 visualizes 10-frame pore-scale facial trajectories from PoreTrack3D, including frontal 2D projections and corresponding 3D trajectories displayed with a colored facial point cloud as a spatial reference. To facilitate visual inspection, trajectories of different types are color-coded: green denotes the primary trajectories that are successfully tracked across time, blue indicates the supplementary trajectories obtained through temporal completion, and red represents supporting trajectories derived from neighboring keypoints. For clarity, only the trajectories over the previous five frames (out of the total 150) are displayed for each frame, and the sequences are arranged from easy to difficult according to the order of dataset construction. The same ordering is consistently applied in all subsequent visualizations and tables.
@@ IV Evaluation on PoreTrack3D
This section evaluates the performance of state-of-the-art dynamic 3D Gaussian splatting methods on the PoreTrack3D benchmark, outlining the experimental setup and reporting both quantitative and qualitative results. The aim is to assess how existing reconstruction paradigms handle fine-scale facial dynamics and to establish a unified baseline for future work on millimeter-level 3D tracking.
@@ IV-B Performance of Baselines
063 Dyn3DGS 0.87 28.49 0.27 0.25 21.16 20.66 0.20 43.71 56.66
Dyn3DGS-robot 0.77 19.25 0.39 0.28 19.08 1.33 0.28 42.40 46.67
4DGS 0.87 31.73 0.17 0.10 81.50 100 0.27 37.23 38.09
GSAvatars 0.98 39.94 0.01 0.12 71.41 100 0.12 71.11 85.33
124 Dyn3DGS 0.87 32.21 0.28 0.13 58.33 60.33 0.20 39.07 52.95
Dyn3DGS-robot 0.78 18.72 0.40 0.25 30.33 6.3 0.24 43.10 48.38
4DGS 0.85 31.44 0.25 0.18 39.58 83.33 0.24 29.25 53.57
GSAvatars 0.97 40.32 0.02 0.08 79.16 96.66 0.16 52.83 79.71
056 Dyn3DGS 0.92 33.17 0.23 0.16 56.16 27.33 0.15 57.50 69.42
Dyn3DGS-robot 0.87 22.00 0.38 2.59 5.16 6.67 3.89 1.59 4.28
4DGS 0.91 35.50 0.19 0.30 24.50 26.00 0.38 11.64 14.28
GSAvatars 0.97 37.33 0.03 0.13 63.16 88.00 0.18 47.28 52.19
038 Dyn3DGS 0.94 33.46 0.17 0.19 46.41 76.33 0.22 44.83 52.85
Dyn3DGS-robot 0.89 19.67 0.28 0.36 20.25 4.66 0.40 8.64 19.23
4DGS 0.94 35.13 0.12 0.12 71.50 81.66 0.48 14.76 33.90
GSAvatars 0.98 39.86 0.01 0.24 26.83 51.33 0.18 47.57 56.38
196 Dyn3DGS 0.90 27.80 0.26 0.37 2.33 2.66 0.21 40.67 53.52
Dyn3DGS-robot 0.84 18.43 0.35 0.27 9.5 24.00 0.38 27.57 33.23
4DGS 0.88 26.05 0.23 0.10 82.16 100 0.82 4.76 1.23
GSAvatars 0.97 37.07 0.02 0.11 74.16 100 0.19 47.57 69.43
033 Dyn3DGS 0.93 33.26 0.24 0.12 75.16 100 0.16 51.14 66.28
Dyn3DGS-robot 0.89 22.75 0.34 0.36 22.83 29.33 0.29 29.85 38.28
4DGS 0.94 36.68 0.13 0.34 12.50 4.00 0.32 19.42 23.23
GSAvatars 0.98 40.20 0.01 0.11 69.50 94.66 0.20 43.71 54.85
264 Dyn3DGS 0.92 31.31 0.21 0.18 48.66 100 0.26 31.90 50.76
Dyn3DGS-robot 0.84 18.79 0.34 0.18 43.50 100 0.63 1.07 4.28
4DGS 0.91 33.52 0.15 0.10 69.83 84.66 0.34 18.85 24.38
GSAvatars 0.97 39.97 0.01 0.06 98.33 100 0.23 34.38 31.61
031 Dyn3DGS 0.96 35.09 0.16 0.17 50.91 48.33 0.37 22.35 22.09
Dyn3DGS-robot 0.93 22.02 0.25 0.72 16.50 2.00 0.64 5.16 11.23
4DGS 0.96 35.54 0.08 0.47 12.41 7.00 0.56 8.00 12.28
GSAvatars 0.98 40.75 0.01 0.13 67.16 100 0.34 18.81 20.85
@@
Dyn3DGS achieves solid overall performance, serving as a strong baseline for dynamic 3D Gaussian optimization. By directly updating Gaussian parameters under physical regularization, it can capture fine-scale motion and produces competitive trajectory accuracy. However, the lack of explicit structural priors makes it more susceptible to local drift, especially in regions with complex, high-frequency facial motion.
Dyn3DGS-robot exhibits the lowest scores among all evaluated baselines. The method models dynamics using a reduced set of control points and enforces smoothly varying deformations across time, which leads to more conservative motion updates. While this formulation is effective for maintaining stable global structure, it tends to underfit the high-frequency, fine-grained displacements present in pore-scale facial motion. As a result, the reconstructed trajectories become overly smoothed, causing reductions in both reconstruction fidelity and point-level tracking accuracy.
4DGS, which models dynamics through a learned continuous deformation field parameterized by a HexPlane encoder, attains reconstruction metrics comparable to Dyn3DGS. Its deformation field produces smooth temporal evolution, which benefits large-scale, low-frequency motions such as landmark trajectories. However, this smoothness makes it less capable of representing sharp, fine-scale pore-level deformations.
@@ V Conclusion
We have introduced PoreTrack3D, the first benchmark for dynamic 3D Gaussian splatting in pore-scale, non-rigid facial trajectory tracking. The dataset contains over 440,000 trajectories, including more than 52,000 tracks that exceed 10 frames and 68 manually verified tracks spanning the entire 150-frame sequence. Such long-range, densely sampled correspondences are rarely available in existing geometric datasets and provide strong temporal constraints for evaluating fine-scale, non-rigid surface motion.
By capturing both macroscopic facial landmarks and microscopic pore-level features, PoreTrack3D bridges the gap between coarse landmark tracking and high-precision surface dynamics. Its data creation pipeline—combining automated extraction with targeted human validation—yields a benchmark that is both scalable and reliable for studying subtle, physically consistent deformation.
Using this benchmark, we establish the first performance baseline for dynamic 3D Gaussian splatting methods, systematically analyzing their ability to model fine-grained, non-rigid motion. Current 3D Geometry Foundation Models have rarely been evaluated on facial, non-rigid, long-sequence scenarios due to the lack of suitable ground-truth correspondences. PoreTrack3D fills this gap by providing the type of long, dense tracks required for such assessments, thereby enabling—rather than requiring—future work to explore GFMs under this challenging regime.
Looking forward, PoreTrack3D advances the quantitative study of pore-scale facial deformation and provides a valuable supervisory signal for future methods across neural deformation fields, Gaussian-based dynamic modeling, 3D scene-flow estimation, and 3D Geometry Foundation Models. We believe this benchmark establishes an important foundation for precise, interpretable, and physically grounded modeling of 3D facial motion.

### v2

@@ III-F PoreTrack3D Dataset
Based on these calibrated data, we reconstruct a 3D point cloud for the initial frame and generate dense temporal trajectories through our proposed four-stage trajectory reconstruction framework. In total, PoreTrack3D contains over 440,000 trajectories, among which more than 52,000 extend beyond 10 frames. Of these, 68 primary trajectories span the full 150 frames and have been manually reviewed, as illustrated in Fig. 5. These trajectories cover both conventional facial landmarks and pore-scale surface keypoints, enabling quantitative analysis of subtle facial deformation. Specifically, among the 68 full-length trajectories, 20 trajectories correspond to facial landmarks and 48 trajectories correspond to pore-scale keypoints. Notably, trajectories longer than 10 frames constitute less than 15% of the total, underscoring the challenge of achieving temporally coherent keypoint tracking under realistic motion and illumination variations.
Fig. 6 visualizes 10-frame pore-scale facial trajectories from PoreTrack3D, including frontal 2D projections and corresponding 3D trajectories displayed with a colored facial point cloud as a spatial reference. To facilitate visual inspection, trajectories of different types are color-coded: green denotes the primary trajectories that are successfully tracked across time, blue indicates the completion-derived trajectories obtained through temporal completion, and red represents supporting trajectories derived from neighboring keypoints. For clarity, only the trajectories over the previous five frames (out of the total 150) are displayed for each frame, and the sequences are arranged from easy to difficult according to the order of dataset construction. The same ordering is consistently applied in all subsequent visualizations and tables.
@@ IV Evaluation on PoreTrack3D
This section evaluates the performance of state-of-the-art dynamic 3D Gaussian splatting methods on the PoreTrack3D benchmark, outlining the experimental setup and reporting both quantitative and qualitative results. The aim is to assess how existing reconstruction paradigms handle subtle facial dynamics and to establish a unified baseline for future work on millimeter-level 3D tracking.
@@ IV-B Performance of Baselines
063 Dyn3DGS 0.87 28.49 0.27 2.56 21.16 20.66 2.03 43.71 56.66
Dyn3DGS-robot 0.77 19.25 0.39 2.81 19.08 1.33 2.81 42.40 46.67
4DGS 0.87 31.73 0.17 1.56 61.06 81.46 2.71 37.23 38.09
GSAvatars 0.98 39.94 0.01 1.26 71.41 98.00 1.78 65.30 74.33
124 Dyn3DGS 0.87 32.21 0.28 1.31 58.33 60.33 2.01 39.07 52.95
Dyn3DGS-robot 0.78 18.72 0.40 2.52 30.33 6.3 2.47 43.10 48.38
4DGS 0.85 31.44 0.25 1.83 39.58 83.33 2.46 29.25 53.57
GSAvatars 0.97 40.32 0.02 0.81 79.16 96.66 1.64 52.83 79.71
056 Dyn3DGS 0.92 33.17 0.23 1.63 54.83 27.33 1.52 57.50 69.42
Dyn3DGS-robot 0.87 22.00 0.38 25.91 5.16 6.67 38.94 1.59 4.28
4DGS 0.91 35.50 0.19 3.02 24.50 26.00 3.82 11.64 14.28
GSAvatars 0.97 37.33 0.03 1.32 63.16 88.00 2.98 28.97 35.22
038 Dyn3DGS 0.94 33.46 0.17 1.92 46.41 76.33 2.27 44.83 52.85
Dyn3DGS-robot 0.89 19.67 0.28 3.65 20.25 4.66 4.08 8.64 19.23
4DGS 0.94 35.13 0.12 1.21 71.50 81.66 4.82 14.76 33.90
GSAvatars 0.98 39.86 0.01 2.44 26.83 51.33 1.99 45.50 68.16
196 Dyn3DGS 0.90 27.80 0.26 3.72 2.33 2.66 2.16 40.67 53.52
Dyn3DGS-robot 0.84 18.43 0.35 2.74 9.5 24.00 3.83 27.57 33.23
4DGS 0.88 26.05 0.23 2.79 41.08 56.66 6.14 3.94 7.77
GSAvatars 0.97 37.07 0.02 1.41 68.91 100 1.92 47.57 69.43
033 Dyn3DGS 0.93 33.26 0.24 2.05 39.55 29.55 1.61 51.14 66.28
Dyn3DGS-robot 0.89 22.75 0.34 3.62 22.83 29.33 2.91 29.85 38.28
4DGS 0.94 36.68 0.13 3.43 12.50 4.00 3.22 19.42 23.23
GSAvatars 0.98 40.20 0.01 2.81 28.49 50.88 2.07 43.71 54.85
264 Dyn3DGS 0.92 31.31 0.21 1.81 48.66 100 2.63 31.90 50.76
Dyn3DGS-robot 0.84 18.79 0.34 1.88 43.50 100 6.31 1.07 4.28
4DGS 0.91 33.52 0.15 1.09 69.83 84.66 3.42 18.85 24.38
GSAvatars 0.97 39.97 0.01 0.85 89.91 100 2.32 34.38 31.61
031 Dyn3DGS 0.96 35.09 0.16 1.72 50.91 48.33 3.78 22.35 22.09
Dyn3DGS-robot 0.93 22.02 0.25 7.28 16.50 2.00 6.42 5.16 11.23
4DGS 0.96 35.54 0.08 4.73 12.41 7.00 5.63 8.00 12.28
GSAvatars 0.98 40.75 0.01 1.58 57.27 57.11 3.59 16.62 16
@@
Dyn3DGS achieves solid overall performance, serving as a strong baseline for dynamic 3D Gaussian optimization. By directly updating Gaussian parameters under physical regularization, it can capture subtle motion and produces competitive trajectory accuracy. However, the lack of explicit structural priors makes it more susceptible to local drift, especially in regions with complex, high-frequency facial motion.
Dyn3DGS-robot exhibits the lowest scores among all evaluated baselines. The method models dynamics using a reduced set of control points and enforces smoothly varying deformations across time, which leads to more conservative motion updates. While this formulation is effective for maintaining stable global structure, it tends to underfit the high-frequency, subtle displacements present in pore-scale facial motion. As a result, the reconstructed trajectories become overly smoothed, causing reductions in both reconstruction fidelity and point-level tracking accuracy.
4DGS, which models dynamics through a learned continuous deformation field parameterized by a HexPlane encoder, attains reconstruction metrics comparable to Dyn3DGS. Its deformation field produces smooth temporal evolution, which benefits large-scale, low-frequency motions such as landmark trajectories. However, this smoothness makes it less capable of representing sharp, pore-scale deformations.
@@ V Conclusion
We have introduced PoreTrack3D, the first benchmark for dynamic 3D Gaussian splatting in pore-scale, non-rigid facial trajectory tracking. The dataset contains over 440,000 trajectories, including more than 52,000 tracks that exceed 10 frames and 68 manually verified tracks spanning the entire 150-frame sequence. Such long-range, densely sampled correspondences are rarely available in existing geometric datasets and provide strong temporal constraints for evaluating subtle, non-rigid surface motion.
By capturing both macroscopic facial landmarks and microscopic pore-scale features, PoreTrack3D bridges the gap between coarse landmark tracking and high-precision surface dynamics. Its data creation pipeline—combining automated extraction with targeted human validation—yields a benchmark that is both scalable and reliable for studying subtle, physically consistent deformation.
Using this benchmark, we establish the first performance baseline for dynamic 3D Gaussian splatting methods, systematically analyzing their ability to model subtle, non-rigid motion. Current 3D Geometry Foundation Models (GFMs) have rarely been evaluated on facial, non-rigid, long-sequence scenarios due to the lack of suitable ground-truth correspondences. PoreTrack3D fills this gap by providing the type of long, dense tracks required for such assessments, thereby enabling—rather than requiring—future work to explore GFMs under this challenging regime.
Looking forward, PoreTrack3D advances the quantitative study of pore-scale facial deformation and provides a valuable supervisory signal for future methods across neural deformation fields, Gaussian-based dynamic modeling, 3D scene-flow estimation, and GFMs. We believe this benchmark establishes an important foundation for precise, interpretable, and physically grounded modeling of 3D facial motion.